# CaloQVAE : Simulating high-energy particle-calorimeter interactions using hybrid quantum-classical generative models

### v4

@@ 4 Performance Evaluation
Figure 2: Comparison of selected shower shape variables for positrons (left sub-panel pair) and charged pions (right sub-panel pair) in a calorimeter. Each pair of sub-panels shows the distributions of two key variables: (1) Fraction of energy: The energy deposited in the first layer compared to the sum of energy in all layers. (2) Sparsity: The ratio of the number of hits in the first layer to the total number of hits in all layers. Each variable is plotted for both particle types to illustrate the close alignment among the results from GEANT4, DVAE (classical) and QVAE (quantum) models.
@@ 5 Conclusion
This work is the first demonstration of the application of a quantum annealing device to the computationally expensive simulation of particle showers at the Large Hadron Collider. We have shown that it is possible to utilize the Quantum Processing Unit (D-Wave Chimera 2000Q) to generate RBM samples which can be used to generate particle showers. The QPU-generated events show good alignment with the classically Monte Carlo-generated events. Despite the sampling rate remaining limited, due to present engineering constraints, the raw QPU annealing rate can be as short as 20 $\mathrm{\mu s}$ per sample, which theoretically allows for a potential of 20-fold speed increase over GPU sample generation. While further work remains in evaluating on more realistic simulation and a wider range of physics topologies, the initial results demonstrate CaloQVAE to be a promising application of quantum computing to open research questions in fundamental physics.

### v5

@@ 4 Performance Evaluation
Figure 2: Comparison of selected shower shape variables for three clusters ($\pi^{+}$, $\gamma$, and $e^{+}$). The rows of sub-panels show the distributions of three key variables: (1) Total Energy: The total energy observed across all layers. (2) Energy Fraction: The fraction of energy deposited in the layer 1 (middle layer) compared to the total energy in all layers. (3) Sparsity: The ratio of the number of hits in the layer 1 to the total number of hits across all layers. Each sub-panel illustrates the close alignment of results from the GEANT4, DVAE (classical), and QVAE (quantum) models.
@@ 5 Conclusion
This work is the first demonstration of the application of a quantum annealing device to the computationally expensive simulation of particle showers at the Large Hadron Collider. We have shown that it is possible to utilize the Quantum Processing Unit (D-Wave Chimera 2000Q) to generate RBM samples which can be used to generate particle showers. The QPU-generated events show good alignment with the classically Monte Carlo-generated events. While the present engineering challenges may not be fully mitigated in the future, the raw QPU annealing rate of 20 $\mu s$ hints at a potential order of magnitude speed increase over GPU sample generation. The initial results demonstrate that CaloQVAE is a promising application of quantum computing to open research questions in fundamental physics. However, further work is needed to evaluate its performance on more realistic simulations, such as the CaloChallenge datasets michele_faucci_giannelli_2023_8099322 ; faucci_giannelli_2022_6366271 ; faucci_giannelli_2022_6366324 and across a wider range of physics topologies.